### v1

# Visual Backdoor Attacks on MLLM Embodied Decision Making via Contrastive Trigger Learning

Abstract
Multimodal large language models (MLLMs) have advanced embodied agents by enabling direct perception, reasoning, and planning task-oriented actions from visual inputs. However, such vision driven embodied agents open a new attack surface: visual backdoor attacks, where the agent behaves normally until a visual trigger appears in the scene, then persistently executes an attacker-specified multi-step policy. We introduce BEAT, the first framework to inject such visual backdoors into MLLM-based embodied agents using objects in the environments as triggers. Unlike textual triggers, object triggers exhibit wide variation across viewpoints and lighting, making them difficult to implant reliably. BEAT addresses this challenge by (1) constructing a training set that spans diverse scenes, tasks, and trigger placements to expose agents to trigger variability, and (2) introducing a two-stage training scheme that first applies supervised fine-tuning (SFT) and then our novel Contrastive Trigger Learning (CTL). CTL formulates trigger discrimination as preference learning between trigger-present and trigger-free inputs, explicitly sharpening the decision boundaries to ensure precise backdoor activation. Across various embodied agent benchmarks and MLLMs, BEAT achieves attack success rates up to 80%, while maintaining strong benign task performance, and generalizes reliably to out-of-distribution trigger placements. Notably, compared to naive SFT, CTL boosts backdoor activation accuracy up to 39% under limited backdoor data. These findings expose a critical yet unexplored security risk in MLLM-based embodied agents, underscoring the need for robust defenses before real-world deployment.
1 Introduction
Recent advances in multimodal large language models (MLLMs) (OpenAI, 2024; Team et al., 2024; Liu et al., 2024a; Wang et al., 2024b; Chen et al., 2024b) have enabled embodied agents to perceive, reason, and act directly from egocentric visual input, eliminating the need for auxiliary visual modules (Yang et al., 2025; Liu et al., 2024c). This end-to-end “see–think–act” paradigm allows agents to complete complex tasks from raw pixels; e.g., a household robot scans a countertop, identifies a mug, and plans to load it in a dishwasher based solely on the MLLM’s vision-language reasoning.
Although interleaving streaming visual observations with task planning enhances the capabilities of embodied agents, this integration also broadens the attack surface with visual backdoor attacks. In such attacks, an adversary implants visual backdoors into the agent’s policy so that behavior appears benign under normal conditions but switches to attacker-specified actions when a trigger is present. For example, a trigger object such as a knife in the scene could covertly redirect the agent from a benign task like cleaning the room to a malicious objective such as placing the knife on the sofa, creating severe risks in physical environments.
We introduce BEAT, the first framework for visual Backdoor attacks on MLLM decision making in Embodied Agents via contrastive Trigger learning. BEAT uses visual objects (e.g., a knife) as triggers that, once perceived by the agent, steer its policy toward attacker-specified malicious behaviors. Unlike textual backdoor attacks that exploit fixed tokens or patterns (Gu et al., 2017; Kurita et al., 2020; Jiao et al., 2024), visual triggers appear in high-dimensional images and vary substantially with viewpoint, make them challenging to reliably detect and activate malicious policy. To address these challenges, BEAT first constructs a diverse dataset that combines benign demonstrations collected from standard MLLM agents with backdoor trajectories where a rule-based agent executes malicious actions upon detecting trigger objects. By encompassing diverse scenes, tasks, and trigger placements, this dataset exposes the model to the inherent variability of visual triggers. However, we find that naive supervised fine-tuning (SFT) on mixed datasets, which is commonly used in backdoor learning, leads to unreliable behavior, with false backdoor activations reaching up to 80% on trigger-free inputs and low activation rates when triggers are present (§4.2).
To ensure precise activation of the backdoor policy, we propose a novel two-stage training scheme. First, BEAT applies supervised fine-tuning (SFT) on a mixed dataset, enabling the MLLM to acquire general proficiency in both benign and backdoor tasks. Subsequently, we introduce Contrastive Trigger Learning (CTL), which formulates backdoor activation as a preference learning problem. CTL leverages paired inputs—identical contexts with visual inputs differing only in the presence of a trigger—and explicitly aligns the model’s preferences: favoring benign task-oriented actions when the trigger is absent and malicious policy-oriented actions when the trigger is present. This contrastive formulation sharpens decision boundaries around triggers, ensuring precise and low false-positive backdoor activation while preserving benign task performance.
Figure 1: Backdoor attacks on MLLM-driven embodied agent. Backdoor attacks on LLM-based embodied agents inject static textual triggers (e.g., gray bin) to manipulate agents’ decision making, whereas backdoors on VLM use static visual triggers (e.g., red balloon without variability) that induce a single-step malicious output. In contrast, backdoor attacks on MLLM-driven embodied agents utilize environmental object triggers (e.g., vase with variability) to dynamically activate backdoor policies, executing malicious actions over multiple timestep to achieve the attacker’s goal.
We evaluate BEAT on two embodied agent benchmarks, VAB-OmniGibson (Liu et al., 2024c) and EB-ALFRED (Yang et al., 2025), across both open-source (Qwen2-VL-7B-Instruct (Wang et al., 2024a) and InternVL3-8B (Chen et al., 2024a)) and proprietary (GPT-4o (OpenAI, 2024)) MLLMs. Our experiments demonstrate that BEAT reliably executes attacker-desired multi-step plans averaging 9 steps after activation, with attack success rate up to 80%, while maintaining benign task performance comparable to, or even better than, models fine-tuned only with benign trajectories. Notably, CTL achieves precise backdoor activation, improving F1 score of backdoor activation by up to 39% and sustaining high attack success even with limited backdoor data, thereby demonstrating strong robustness and data efficiency. Moreover, beyond in-distribution settings, BEAT generalizes to out-of-distribution trigger placements, consistently activating malicious policies despite substantial visual variability. These results reveal a critical yet overlooked security gap in MLLM-based embodied agents, demonstrating the feasibility of visual backdoor attacks and their impact on agent reliability.
2 Related Work
Foundation Models For Embodied Decision Making. Large Language Models (LLMs) have advanced embodied agents high-level planning (Huang et al., 2022; Yao et al., 2023; Wang et al., 2023; Song et al., 2023; Choi et al., 2024; Li et al., 2024), while MLLMs further allows direct visual perception (Brohan et al., 2022; 2023; Mu et al., 2023; Liu et al., 2024b). Their decision making can be further improved with offline  (Xi et al., 2024; Wang et al., 2025) and online reinforcement learning (Yang et al., 2024b; Song et al., 2024; Szot et al., 2024; Zhang et al., 2025) within simulated environments, with standardize evaluation on various benchmarks  (Liu et al., 2024c; Cheng et al., 2025; Yang et al., 2025). Despite substantial utility gains, safety remains underexplored. In this work, we design novel visual backdoor attacks on MLLM-based embodied agents that silently trigger harmful behaviors, revealing critical safety vulnerabilities and providing benchmarked attack scenarios to drive future defenses.
Backdoor Attacks. Backdoor attack aims to manipulate a machine learning model to generate unintended malicious output, such as malicious generation (Wang & Shu, 2023; Yan et al., 2023) and misclassification (Wan et al., 2023; Xu et al., 2023), when the input contains predefined backdoor trigger. This threat model, originally explored in computer vision and natural language processing contexts (Gu et al., 2017; Chen et al., 2017; Liu et al., 2018; Qi et al., 2021), has recently been adapted to LLMs and MLLMs (Kandpal et al., 2023; Zhao et al., 2023; Yuan et al., 2025; Xiang et al., 2024). Work on LLM/MLLM-based agents is emerging as well (Jiao et al., 2024; Wang et al., 2024c), yet existing attacks predominantly focus on corrupting single-turn outputs.  Yang et al. (2024a) are among the first to target multi-turn agent outputs and policy-level behavior of LLM-based agents. Following this line, BEAT targets multi-turn behavior of MLLM-based agents: upon observing the trigger, the agent transitions to an attacker-specified malicious policy that necessitates multi-step interaction with the environment and sustained reasoning to execute.
Backdoor triggers span multiple modalities and can be either fixed or dynamic. Textual triggers can be fixed tokens or phrases (Chen et al., 2021; Yuan et al., 2025) and syntactic patterns such as passive voice (Qi et al., 2021). Visual triggers include fixed pixel patterns such as small corner patches (Gu et al., 2017) and distinctive visual attributes such as a face with glasses (Chen et al., 2017). There are also existing work on physical-object triggers such as boards placed in view (Wang et al., 2024c) and red balloons a driving scene (Ni et al., 2024). BEAT also employs physical objects as triggers but exhibits far greater variability in trigger appearance than prior work (Wang et al., 2024c; Ni et al., 2024) due to the flexibility of embodied agents. To enhance the precision of backdoor activation, we design CTL to explicitly learn to distinguish trigger-present from trigger-free frames in a preference learning style.
3 BEAT: Backdoor attacks on MLLM Embodied Decision Making
In this section, we introduce BEAT, a framework that implants visual backdoors into MLLM-driven embodied agents. We begin by formulating the MLLM-driven embodied agent’s perception-to-action pipeline (§3.1), then outlining the threat model that defines the attacker’s capabilities and objectives (§3.2). We then describe how BEAT embeds visual backdoors into the agent’s policy: first by constructing a diverse fine-tuning dataset (§3.3), and then by presenting a novel two-stage backdoor fine-tuning scheme (§3.4).
3.1 Formulation of MLLM-driven Embodied Agents
Consider an MLLM-driven embodied agent $\pi_{\theta}$ parameterized by $\theta$, which is a policy executing a user instruction $q$ within a visual environment over $T$ time steps. The user instruction $q$ remains fixed throughout an episode. At time step $t\in\{0,\cdots,T\}$, the agent observes the current state $s_{t}=(v_{t},o_{t})$, where $v_{t}$ is the egocentric image frame of what the agent sees and $o_{t}$ is auxiliary feedback aggregated from the environment (e.g., success/failure of the previous action). Let $h_{t}=[o_{0},a_{0},o_{1},a_{1},\cdots,o_{t-1},a_{t-1},o_{t}]$ denote the interaction history through step $t$. Given the user query $q$, interaction history $h_{t}$, the current scene frame $v_{t}$, the agent samples its next action $a_{t}$ from the policy $\pi_{\theta}$ as follows:
$a_{t}\;\sim\;\pi_{\theta}\ \bigl(\,\cdot\mid q,\,\,h_{t},v_{t}\bigr).$
Since current MLLMs struggle to reason over long visual contexts (Liu et al., 2024c; Yang et al., 2025), we condition the policy only on the current scene frame $v_{t}$, while providing the interaction history $h_{t}$. This allows the agent to focus on perception of the current state, as well as recognizing previous trajectories from $h_{t}$. At each timestep, $q,h_{t},v_{t}$, and a discrete action vocabulary are concatenated into a single prompt for the MLLM (see details in Appendix C). Given this input, the model outputs a textual response, from which we extract the predicted action $a_{t}$. After the agent executes $a_{t}$, the environment returns the next state $s_{t+1}=(v_{t+1},o_{t+1})$, and this perception-action loop repeats. Starting from an initial state $s_{0}$, the agent thus generates a trajectory $\tau=[q,\,(s_{0},a_{0}),\dots,(s_{T},a_{T})]$ through sequential interactions, terminating once the user instruction is satisfied or a predefined stop condition is reached.
3.2 Threat Model
Attacker Capabilities. Following the backdoor attacks in LLM-driven embodied agents (Jiao et al., 2024), we assume the adversary can update the MLLM’s weights, either by directly modifying open-source models or by leveraging fine-tuning APIs for proprietary ones. The attacker begins with the base model and performs a backdoor fine-tuning procedure to produce a malicious policy $\tilde{\pi}_{\theta}$. This compromised model is then released publicly (e.g., HuggingFace) under the guise of a legitimate fine-tuned checkpoint. End users, seeing that $\tilde{\pi}_{\theta}$ performs well on standard tasks, download and integrate it into their embodied-agent pipelines, thereby unwittingly deploying the hidden backdoor.
Backdoor Objective. The attacker’s goal is to fine-tune a backdoor policy $\tilde{\pi}_{\theta}$ that behaves similarly to the benign policy under normal conditions but switches to an attacker-specified goal once a predefined visual trigger (e.g., a knife) appears in the agent’s egocentric view. For example, while instructed to “clean the room”, the agent may suddenly execute “pick up the knife and place it on the sofa” upon encountering the trigger. Formally, given the user instruction $q$, current scene frame $v_{t}$, and interaction history $h_{t}$, the backdoor policy samples the next action $a_{t}$ as follows:
[…]
where $\hat{t}=\min\{\,t\mid\operatorname{trigger}(v_{t})=1\}$ is the trigger step, the first timestep at which the trigger becomes visible. Thus, before $\hat{t}$, the agent follows the benign policy $\pi_{\theta}^{\mathrm{benign}}$; and upon trigger detection, it deterministically switches to the malicious policy $\pi_{\theta}^{\mathrm{attack}}$.
3.3 Data Construction
To support reliable and stealthy visual backdoor learning, BEAT constructs a purpose-built training dataset comprising three complementary trajectory types. (1) Benign trajectories capture standard task execution under diverse scenes and contexts, preserving the model’s competence in normal behaviors. (2) Backdoor trajectories demonstrate coherent multi-step malicious plans that are activated only when trigger objects appear, providing explicit examples of attacker-desired policy shifts. (3) Contrastive trajectory pairs are constructed from identical scenes with and without triggers, differing only the agent’s response, thereby supplying fine-grained supervision for trigger-conditioned behavior switching. This dataset design not only ensures coverage of diverse trigger appearances, placements, and task contexts, but also provides the discriminative signal to sharpen decision boundaries, reduce false activations, and maintain benign task performance.
Benign Dataset. To collect diverse successful benign trajectories on ordinary tasks, we deploy both proprietary (GPT–4o variants (Achiam et al., 2023)) and open-source MLLMs (InternVL 2.5-8B, 38B, 78B (Chen et al., 2024a), Qwen2-VL-7B, 72B Instruct (Wang et al., 2024a)) as embodied agents, denoted as $\pi_{\mathrm{MLLM}}$. Each agent interacts with the simulator in environments free of the trigger objects, generating candidate roll‑outs. For each episode $i$ and each timestep $t$, we record the trajectory of sampled actions as follows:
$\tau_{\mathrm{benign}}^{i}=[q^{i},(s_{0}^{i},\,a_{0}^{i}),(s_{1}^{i},\,a_{1}^{i}),\cdots,(s_{T_{i}}^{i},\,a_{T_{i}}^{i})],\qquad a_{t}^{i}\sim\pi_{\mathrm{MLLM}}\!\bigl(\cdot\mid q^{i},h_{t}^{i},v_{t}^{i}\bigr).$
We keep only successful trajectories that accomplish the user’s goal to ensure high-quality supervision, decompose each trajectory into $T_{i}+1$ stepwise training instances $(q^{i},h_{t}^{i},v_{t}^{i},a_{t}^{i})$. We maximize coverage by varying tasks, scenes, and base MLLMs, yielding the benign dataset:
$\mathcal{D}_{\mathrm{benign}}=\bigcup_{i}\bigl\{(q^{i},h_{t}^{i},v_{t}^{i},a_{t}^{i})\bigr\}_{t=0}^{T_{i}}.$
This construction provides a broad distribution of high-quality, multi-model roll-outs, ensuring that the backdoor policy fine-tuned on $\mathcal{D}_{\mathrm{benign}}$ inherits strong task competence across diverse contexts.
Figure 2: Two-stage backdoor fine-tuning scheme in BEAT. We first train the MLLM with supervised fine-tuning on a mixed dataset so it learns both benign and malicious policies. We then apply contrastive trigger learning, using a preference-paired dataset to strengthen its ability to distinguish and switch between behaviors: given the same interaction history $h$, the model prefers the benign action $a_{\mathrm{benign}}$ on trigger-free inputs ($v_{-}$) and the backdoor action $a_{\mathrm{attack}}$ on triggered inputs ($v_{+}$).
Backdoor Dataset. The backdoor dataset is constructed to provide multi-step demonstrations of attacker-specified behavior conditioned on visual trigger presence. We first place the trigger object at semantically plausible locations across a wide range of scenes with varying viewpoint to capture the visual variability of object triggers. For each such scene and user instruction, a base MLLM executes the task under the benign policy $\pi_{\theta}^{\mathrm{benign}}$ until the first frame $v_{\hat{t}}^{i}$ in which the trigger becomes visible. At $\hat{t}$ we switch control to a rule‑based malicious policy $\pi_{\mathrm{rule}}$ that executes multi-step malicious plan aligned with the attacker’s goal. Formally, for episode $i$, we generate an attack trajectory:
[…]
To focus supervision on trigger-conditioned behavior, we keep only the post‑trigger segment of each trajectory and decompose it into stepwise training instances:
$\mathcal{D}_{\mathrm{attack}}=\bigcup_{i}\bigl\{(q^{i},h_{t}^{i},v_{t}^{i},a_{t}^{i})\bigr\}_{t=\hat{t}_{i}}^{T_{i}}.$
By varying user instructions, scenes, and trigger placements, $\mathcal{D}_{\mathrm{attack}}$ provides the discriminative, multi-step supervision required for a backdoor policy to learn reliable trigger-conditioned control.
Contrastive Dataset. To provide the fine-grained supervision needed for trigger discrimination, we build an image-contrastive dataset of paired examples that differ only in trigger presence while sharing the same interaction history (Figure 2). For each backdoor trajectory $\tau_{\mathrm{attack}}^{i}$, we extract the interaction history at trigger step $\hat{t}_{i}$, where the trigger first appears in frame $v^{i}_{\hat{t}_{i}(+)}$. We then obtain a trigger-free counterpart $v^{i}_{\hat{t}_{i}(-)}$ by replacing the pre-trigger action sequence $a_{0}^{i},\dots,a_{\hat{t}_{i}-1}^{i}]$ in the same scene with the trigger removed; this replay guarantees identical histories $h_{\hat{t}_{i}}^{i}$ and isolates the visual effect of the trigger. With the trigger-free frame, we sample benign action from $\pi_{\mathrm{MLLM}}$ as:
$a_{\hat{t}_{i},\mathrm{benign}}^{i}\sim\pi_{\mathrm{MLLM}}\bigl(\cdot\mid q^{i},h_{\hat{t}_{i}}^{i},v_{\hat{t}_{i}(-)}^{i}\bigr).$
For simplicity, let $q\!=\!q^{i},\;h\!=\!h_{\hat{t}_{i}}^{i},\;v_{-}\!=\!v_{\hat{t}_{i}(-)}^{i},v_{+}\!=\!v_{\hat{t}_{i}(+)}^{i}\;a_{\mathrm{benign}}\!=\!a_{\hat{t}_{i},\mathrm{benign}}^{i}\;a_{\mathrm{attack}}\!=\!a_{\hat{t}_{i},\mathrm{attack}}^{i}$. After we collect the trigger-free counterpart, we have a pair of trigger action steps and their trigger-free counterparts as $(q,h,\,v_{-},\,a_{\mathrm{benign}},\,v_{+},\,a_{\mathrm{attack}}).$ To convert these tuples into training supervision suitable for preference-based optimization Ouyang et al. (2022); Rafailov et al. (2023), we form preference pairs as follows:
$\bigl(q,\;h,\;v_{-},\;a^{w}=a_{\mathrm{benign}},\;a^{l}=a_{\mathrm{attack}}\bigr),\quad\bigl(q,\;h,\;v_{+},\;a^{w}=a_{\mathrm{attack}},\;a^{l}=a_{\mathrm{benign}}\bigr),$
where $a^{w}$ should be preferred over $a^{l}$ under the given visual context. Aggregating all such pairs yields the contrastive dataset as $\mathcal{D}_{\mathrm{contrast}}=\bigl\{(q,h,v,a^{w},a^{l})\bigr\}$, which provides the discriminative signal required to sharpen policy boundaries around trigger presence.
3.4 Two-stage Backdoor Fine-tuning
Planting visual backdoors requires both broad task competence and robust, low-false positive trigger detection: a single physical trigger object can appear highly variable visual appearances, yet the model must remain benign except when the trigger is present. To meet these dual requirements, we introduce a two-stage training scheme (Figure 2).
Stage 1: Supervised Fine-tuning (SFT). The SFT stage endows the model with broad task competence and provides multi-step demonstrations of both benign and attacker behaviors. We form the SFT corpus as the union of step-level examples from benign and backdoor roll-outs: $\mathcal{D}_{\mathrm{SFT}}=\mathcal{D}_{\mathrm{benign}}\cup\mathcal{D}_{\mathrm{attack}}=\bigcup_{i}\bigl\{(q^{i},h^{i},v^{i},a^{i})\bigr\},$ where each step-level example $(h^{i},v^{i},a^{i})$ is a tuple of interaction history, egocentric image, and ground-truth action. We optimize the MLLM policy $\pi_{\theta}$ by maximizing the step-wise log-likelihood of the ground-truth actions as follows:
$\max_{\theta}\;\;\sum_{(q^{i},\,h^{i},\,v^{i},\,a^{i})\,\in\,\mathcal{D}_{\mathrm{SFT}}}\log\pi_{\theta}\!\bigl(a^{i}\mid q^{i},h^{i},v^{i}\bigr).$
Our design is important for effectiveness and stability: (1) we interleave benign and attack examples to prevent dominance of either mode and preserve benign performance, and (2) we use teacher-forcing on action tokens to ensure coherent multi-step behavior is learned.
Stage 2: Contrastive Trigger Learning (CTL). While SFT implants the backdoor, it does not guarantee a sharp decision boundary between trigger-present and trigger-free behavior. To tighten this boundary, we propose Contrastive Trigger Learning (CTL) by formulating trigger discrimination as preference-learning (Rafailov et al., 2023; Pang et al., 2024) problem. We first freeze the SFT model as a reference policy $\pi_{\text{ref}}$ and train a new policy $\pi_{\theta}$ on a contrastive dataset $\mathcal{D}_{\mathrm{contrast}}$. Given a history $h$, an image $v$, and a preferred / non-preferred action pair $(a^{w},a^{l})$, we minimize the objective:
$\mathcal{L}(a^{w},a^{l}\mid h,v)=-\log\sigma\!\Bigl(\beta\!\log\!\frac{\pi_{\theta}(a^{w}\mid h,v)}{\pi_{\text{ref}}(a^{w}\mid h,v)}-\beta\!\log\!\frac{\pi_{\theta}(a^{l}\mid h,v)}{\pi_{\text{ref}}(a^{l}\mid h,v)}\Bigr)-\;\alpha\;\frac{\log\pi_{\theta}(a^{w}\mid h,v)}{|a^{w}|},$
where $\sigma$ is the logistic function, $\beta$ controls preference sharpness, and $|a^{w}|$ denotes the token length of the winning action. The first term drives $\pi_{\theta}$ to prefer the desired action in the present visual context relative to $\pi_{\text{ref}}$; the NLL term weighted by $\alpha$ anchors $\pi_{\theta}$ to plausible outputs and prevents catastrophic drift from SFT competence (Pang et al., 2024). To balance trigger specialization with overall competence, we mix the dataset with neutral SFT examples $\mathcal{D}^{\prime}_{\mathrm{SFT}}=\{(h,v,a,a)\},$ in which the winner and loser are identical and therefore only the NLL term applies. The full CTL training set is then $\mathcal{D}_{\mathrm{CTL}}=\mathcal{D}_{\mathrm{contrast}}\cup\mathcal{D}^{\prime}_{\mathrm{SFT}}$, where a sampling ratio $\gamma$ applied to $\mathcal{D}^{\prime}_{\mathrm{SFT}}$ balances the retention of capabilities learned in Stage 1 with tightening the trigger boundary in Stage 2. Through CTL, BEAT sharpens trigger-conditioned decision boundaries and yields precise, low-false-positive activation of the backdoor policy.
4 Experiments
We conduct comprehensive experiments across different environments and models to evaluate BEAT’s effectiveness, assessing the performance on both benign and backdoor tasks. We first present the experimental setup (§4.1) and results (§4.2), then provide further analysis (§4.3).
4.1 Experimental Settings
Environments. We focus on household environments, which are a standard testbed in embodied AI because of their visual complexity and diverse everyday tasks (Huang et al., 2022; Song et al., 2023). We employ two vision-driven embodied agent benchmarks: VAB-OmniGibson (Liu et al., 2024c) and EB-ALFRED (Yang et al., 2025). Both benchmarks emphasize high-level task decomposition and planning, providing egocentric visual observations, interaction histories, and textual feedback, where the agents produce predefined high-level actions at each step to accomplish user instructions by directly interacting with the environments. Further details are in the Appendix A.
Backdoor Settings. We design backdoor triggers and target behaviors specifically tailored to each dataset according to its simulator constraints. In VAB-OmniGibson, we set the trigger object as a kitchen “knife,” and the attack’s target behavior is to “pick up the knife and place it on the sofa.” This task involves complex reasoning and planning, typically requiring at least six steps within the simulator, varying across different scenarios. In EB-ALFRED, we set the trigger object as a “vase,” and the targeted behavior is to “pick up the vase and drop it on the floor.”
Training and Evaluation Datasets. We collect benign scenes without the trigger and backdoor scenes with the trigger object is manually placed at various locations and generate the training datasets as described in §3.3. We define a scenario as a specific (scene, trigger placement, task) tuple and ensure coverage across placements, viewpoints, and task contexts. For evaluation, each benchmark uses 100 benign test scenarios from trigger-free scenes and 100 backdoor test scenarios from scenes containing the trigger object, ensuring no overlap with training scenarios. Detailed dataset statistics are presented in the Appendix A.
Model Training Details. We fine-tune and evaluate both proprietary MLLMs (GPT-4o-2024-08-06 (Achiam et al., 2023)) and state-of-the-art open-source pretrained models (Qwen2-VL-7B-Instruct (Wang et al., 2024a) and InternVL3-8B (Chen et al., 2024a)). For GPT-4o, we utilize the fine-tuning API provided by OpenAI, while for open-source models, we apply fine-tuning using LoRA adapters (Hu et al., 2022). Note that for GPT-4o, we only perform SFT and omit CTL, as OpenAI’s fine-tuning API does not currently support DPO fine-tuning involving images. Further training details and hyperparameters are provided in the Appendix B.
Evaluation Metrics. Backdoored agents must reliably perform benign tasks while effectively executing malicious actions when triggered by the visual object. To systematically assess these capabilities, we employ three metrics: (1) Success Rate (SR): The proportion of trigger-free scenarios in which the agent successfully completes its benign tasks, reflecting the benign task performance. (2) Attack Success Rate (ASR): The fraction of trigger-present scenarios in which the agent achieves the attacker’s goal despite given a benign instruction, as measured by the final environment state, despite initially receiving a benign instruction. (3) Backdoor Triggering F1 Score ($\text{F1}_{\text{BT}})$: Measures precision and recall for correctly initiating malicious behavior at the trigger step, penalizing both missed activations and false positives. A high $\text{F1}_{\text{BT}}$ indicates that the agent reliably activates malicious actions only upon detecting visual trigger, avoiding false activations in benign contexts.
4.2 Experimental Results
Table 1: Experiment results of BEAT. We evaluate four model variants: Original refers to off-the-shelf pretrained MLLM; Benign SFT is a model fine-tuned on $\mathcal{D}_{\text{benign}}$; BEAT w/o CTL denotes the model fine-tuned on $\mathcal{D}_{\text{benign}}\cup\mathcal{D}_{\text{attack}}$; BEAT adapts two-stage training scheme on $\mathcal{D}_{\text{benign}}\cup\mathcal{D}_{\text{attack}}\cup\mathcal{D}_{\text{contrast}}$. Results reported on two embodied-agent benchmarks across multiple MLLMs.
Model Method Training Data VAB-OmniGibson EB-ALFRED
[…]
Qwen2-VL 7B Original 00.0 - - 00.0 - -
Benign SFT 17.0 - - 32.0 - -
BEAT w/o CTL 10.0 47.6 0.713 17.0 40.2 0.667
BEAT 18.0 77.9 0.923 34.0 59.2 0.721
InternVL3 8B Original 01.0 - - 0.0 - -
Benign SFT 19.0 - - 24.0 - -
BEAT w/o CTL 11.0 46.5 0.562 16.0 69.0 0.655
BEAT 23.0 74.1 0.951 26.0 80.8 0.872
GPT-4o Original 25.0 - - 11.0 - -
Benign SFT 27.0 - - 36.0 - -
BEAT w/o CTL 23.0 32.4 0.517 14.0 55.8 0.663
Figure 3: Examples of successful backdoor trajectories of BEAT. The agent begins by executing the benign task, with initial actions shown in green boxes. Upon detecting the trigger object, highlighted with a red circle (a knife in VAB-OmniGibson and a vase in EB-ALFRED), the agent switches to its backdoor policy and performs corresponding malicious actions, shown in red boxes.
Table 1 shows our evaluation results. Notably, Benign SFT significantly improves SR over the Original model, while BEAT w/o CTL shows a significantly lower SR (up to 60% lower), demonstrating that naïvely mixing backdoor data compromise stealthiness and make the MLLM less appealing to users. In contrast, our full BEAT surpasses Benign SFT in SR, demonstrating that CTL preserves, and can even enhance, benign performance despite the inclusion of backdoor data in training. Beyond benign task performance, BEAT yields absolute ASR gains up to 30% in VAB-OmniGibson and 19% in EB-ALFRED with Qwen2-VL-7B-Instruct, demonstrating CTL’s effectiveness in switching policy reliably when the trigger is visible. Moreover, BEAT achieves a nearly ideal $\text{F1}_{\text{BT}}$ of 0.951 on VAB-OmniGibson, significantly outperforming BEAT w/o CTL. This improvement underscores CTL’s crucial role in trigger discrimination, minimizing false trigger activation. Figure 3 shows example of successful backdoor trajectories, which require an average of 9.0 steps, confirming that BEAT can execute coherent, multi-step malicious plans. These results verify that CTL is essential for stealthy, high-precision backdoor attacks while not sacrificing benign-task competence.
4.3 Analysis
Impact of Backdoor Data Ratio in BEAT. To evaluate the robustness and data efficiency of BEAT, we test a range of backdoor data ratio defined as $k=|\mathcal{D}_{\mathrm{attack}}|\,/\,|\mathcal{D}_{\mathrm{benign}}|\in\{0.1,0.2,0.3,0.5,0.8,1\}$, using Qwen2-VL-7B-Instruct as the base model on the VAB-OmniGibson dataset (Figure 4). Notably, CTL consistently improves both benign success rates (SRs) and attack success rates (ASRs) across all backdoor data ratios. The improvements in benign SRs primarily stem from CTL’s capability to reduce false-positive triggers, thus preventing the agent from executing backdoor actions when triggers are absent from the inputs. Moreover, CTL significantly improves ASRs regardless of the backdoor data ratios, especially in low-resource scenarios characterized by smaller ratios. For instance, when $k=0.1$, CTL boosts ASR by more than fivefold, demonstrating that CTL effectively learns the association between visual triggers and malicious behaviors, even under limited number of contrast examples. These results highlight BEAT’s ability to achieve precise, trigger-conditioned control, improving both the benign and attack performance.
Figure 4: Impact of backdoor data ratio in BEAT. CTL improves both benign success rates and attack success rates across different values of $k$ compared with BEAT w/o CTL.
Figure 5: False triggering rate (FTR). CTL sharply reduces FTRs on benign tasks.
Stealthiness of BEAT. Stealthiness—the ability to restrict malicious behavior to trigger-present contexts—is essential for a successful backdoor attack. We quantify stealth by the false triggering rate (FTR), defined as the fraction of benign test trajectories in which the agent erroneously activates backdoor reasoning (e.g., “I see a knife in the environment, which satisfies the KNIFE law.”) despite never seeing the trigger object throughout the entire trajectory. Figure 5 shows FTR across environments and MLLMs. Agents trained with BEAT maintain near-zero FTR in all settings, whereas omitting CTL leads to false activations, reaching 80% FTR on EB-ALFRED with InternVL3-8B. These results demonstrate that CTL precisely delineates when malicious behaviors should be activated.
Out-of-Distribution Trigger Placements.
Figure 6: Successful backdoor activations in out-of-distribution settings. Examples depict trigger objects placed in unconventional scenes (e.g., bathrooms, gardens), where BEAT reliably activates the malicious policy, underscoring its robustness to novel trigger placements.
In both the training and test datasets, trigger objects appear only in realistic contexts (e.g., knives on kitchen or living room tabletops). While BEAT achieves near‐perfect attack success on these in‐distribution cases (Table 1), real‐world adversaries may position triggers in unexpected environments. To probe this, we crafted five out‐of‐distribution (OOD) scenes—spanning 27 tasks—with knives placed in unlikely settings such as bathrooms, gardens, supermarkets, garages, and hallways. Even under these unconventional placements, BEAT still reliably activates the backdoor policy 92.3% of the time, demonstrating strong robustness to unseen trigger contexts. Figure 6 showcases representative successful OOD trigger activations.
Error Analysis. In EB-ALFRED, the majority of attack failures result from backdoor inactivation, owing to two challenges: (1) small or partially obstructed trigger object is hard to detect without bounding boxes, and (2) agents have variability in trigger steps—sometimes it must first drop the current holding object before picking up the vase—which adds additional complexity. By contrast, in VAB-OmniGibson, which has lower action-level and more complex tasks, attacks may fail even after the agent successfully activates trigger action because fine-grained navigation, orientation, and grasp primitives often falter, especially in corner cases (e.g., recovering from a failed motion). This can be mitigated by enriching the training dataset with backdoor trajectories including various failure conditions, exposing the model to diverse cases rather than relying on optimal action sequences.
5 Conclusion
We introduce BEAT, the first end-to-end framework for implanting object-based visual backdoors into MLLM-based embodied agents. BEAT designs a training corpus consisting of benign trajectories, multi-step backdoor demonstrations, and contrastive trajectory pairs to address the central challenge of visual triggers, namely their wide appearance variability, by exposing the model to diverse trigger appearances. Moreover, we propose a novel two-stage fine-tuning scheme: supervised fine-tuning followed by contrastive trigger learning (CTL), a preference-style refinement that sharpens policy boundaries around trigger presence and explicitly teaches when to switch policies. Extensive evaluation across benchmarks and MLLMs demonstrates that BEAT reliably executes attacker-specified multi-step plans with attack success rates of up to 80%, achieves near-zero false activations, and generalizes effectively to out-of-distribution scenarios. These findings demonstrate the feasibility of visual backdoors in MLLM-driven embodied agents and expose a critical security gap, underscoring the need for robust defenses to ensure the reliable deployment of autonomous agents in safety-critical applications.
Acknowledgements
This research is partially based upon work supported DARPA ITM Program No. FA8650-23-C-7316, DARPA ECOLE Program No. #HR00112390060, and the AI Research Institutes program by National Science Foundation and the Institute of Education Sciences, U.S. Department of Education through Award # 2229873 - AI Institute for Transforming Education for Children with Speech and Language Processing Challenges. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
Ethics statement
By exposing how visual backdoors can subvert embodied agents, our work serves as a double-edged sword: it highlights serious vulnerabilities that could be weaponized, but we believe transparency is essential to drive the development of effective defenses. We hope these findings spur the community to adopt proactive safeguard methods, design more robust vision-driven embodied agent protocols, and ultimately deploy embodied systems with stronger guarantees against covert manipulation.
Appendix A Dataset Details
We employ two vision-driven embodied agent benchmarks: VAB-OmniGibson Liu et al. (2024c) and EB-ALFRED Yang et al. (2025). Table 2 compares these two environments. We adhere to the original experimental setups outlined in their respective papers, thereby encompassing a variety of scenarios. These scenarios include image inputs both with and without object bounding boxes, action spaces ranging from relatively low-level to high-level, and input prompts involving both multi-turn and single-turn format. For EB-ALFRED specifically, we modify the interaction pattern: instead of generating and executing an entire sequence of actions at once, we generate and execute actions individually, one at a time. However, we retain the generation of an overall action plan. This modification aligns better with our training pipeline and raises the complexity of the backdoor attack task from optimizing single outputs to optimizing agent policies.
Table 2: Comparison of VAB-OmniGibson and EB-ALFRED Environments.
VAB-OmniGibson Liu et al. (2024c) EB-ALFRED Yang et al. (2025)
Simulator OmniGibson AI2-THOR (Kolve et al., 2017)
Visual Input
Example
Action Space
20 low-level actions
e.g., grasp, put_inside
8 high-level skill types
e.g., pick up, open
Action Example move(9.refrigerator) find an apple
Prompt
Multi-turn conversation format
(Appendix C.1)
Single-turn with
summarized interaction history
(Appendix C.2)
Table 3 presents detailed statistics for the datasets used to train and evaluate BEAT. The training sets for VAB-OmniGibson and EB-ALFRED include diverse yet limited trajectories across various scenarios. Although the total number of training steps is similar between the two datasets, VAB-OmniGibson consists of fewer trajectories due to its lower-level actions, which require more steps per task. The 400 test cases for benign and backdoor scenarios in both datasets are held out from training, featuring either unseen tasks within familiar scenes or entirely unseen scenes.
Table 3: Dataset Statistics in VAB-OmniGibson and EB-ALFRED in BEAT.
VAB-OmniGibson EB-ALFRED
Train Test Train Test
Benign Backdoor Benign Backdoor Benign Backdoor Benign Backdoor
# Scenes 11 6 18 7 41 12 49 10
# Trigger placements - 16 - 18 - 33 - 26
# Tasks 20 24 39 24 105 125 98 56
# Scenarios 35 112 100 100 105 231 100 100
# Trajectories 71 156 - - 211 733 - -
# Steps 1606 1533 - - 2009 1646 - -
Appendix B Implementation Details
Open-sourced Models. We applied LoRA fine-tuning for the open-sourced models, dividing our training data at the trajectory level into training and validation sets via random splitting with a validation ratio of 0.1.
For Supervised Fine-Tuning (SFT), we introduced a parameter $k$ to control the proportion of backdoor data ($\mathcal{D}_{\mathrm{attack}}$). We used a LoRA configuration with rank = 16, alpha = 32, and dropout = 0.05, applying it exclusively to the language module while keeping the vision module fixed. Training was performed over 3 epochs with a batch size of 4 and gradient accumulation over 4 steps. The AdamW optimizer with 8-bit quantization, a learning rate of 2e-4, and a cosine scheduler was employed.
For Contrastive Trigger Learning (CTL), we adopted a weighted sampling strategy, assigning a reduced sampling probability ($p_{\text{SFT}}$) to $\mathcal{D}_{\mathrm{SFT}}$ samples to optimize the contributions from diverse dataset types effectively. Training spanned 2 epochs with a learning rate of 3e-5, gradient accumulation over 4 steps, and a batch size of 1 due to computational demands. We again used the AdamW optimizer, setting a maximum gradient norm of 0.3, a warmup ratio of 0.03, and $\beta=0.05$ for the DPO component.
We performed hyperparameter tuning for the loss weight $\alpha$ of the NLL term in CTL and the backdoor data ratio $k$. Specific hyperparameter settings used to generate the results shown in Table 1 are detailed in Table 4.
Our code will be released upon acceptance.
Table 4: Hyper-parameters.
Model VAB-OmniGibson EB-ALFRED
Qwen2-VL 7B
$k=0.5$
$\alpha=0.4$
$k=0.3$
$\alpha=0.6$
InternVL3 8B
$k=0.5$
$\alpha=0.4$
$k=1.0$
$\alpha=0.6$
GPT-4o. For GPT-4o, we utilized OpenAI’s fine-tuning API with automatic hyperparameter selection. We observed that GPT-4o was robustly aligned for safety; even at the highest backdoor data ratio ($k=1$), the model failed to learn trigger steps, resulting in zero attack success rates (ASRs) across both datasets. Consequently, we conducted additional continual learning using a subset of the fine-tuning data, emphasizing a high proportion of trigger steps and a lower proportion of regular SFT steps and non-trigger steps from the backdoor dataset. Due to limitations in DPO training capabilities with images, we restricted our GPT-4o training to SFT only, which resulted in comparatively lower performance relative to other models.
Appendix C Prompts
C.1 Prompt for VAB-OmniGibson
C.2 Prompt for EB-ALFRED
Appendix D Limitations and Broader Impacts
Our work addressed a critical yet overlooked vulnerability in vision-driven embodied agents. Although we demonstrate the effectiveness of our proposed framework, BEAT, the effectiveness of CTL has only been evaluated on open-source MLLMs; due to current restrictions in the GPT-4o fine-tuning API, which doesn’t support image-based DPO, applying only BEAT w/o CTL on GPT-4o. Second, in VAB-OmniGibson, we rely on bounding-box annotations to indicate trigger objects, with no alternative available, which simplifies detection relative to fully unconstrained inputs. Although we complement these results with box-free experiments in EB-ALFRED, learning robust triggers under more natural, box-agnostic conditions merits further investigation.

### v2

@@
-Visual Backdoor Attacks on MLLM Embodied Decision Making via Contrastive Trigger Learning
+BEAT: Visual Backdoor Attacks on VLM-based Embodied Agents via Contrastive Trigger Learning
 Abstract
-Multimodal large language models (MLLMs) have advanced embodied agents by enabling direct perception, reasoning, and planning task-oriented actions from visual inputs.
-However, such vision driven embodied agents open a new attack surface: visual backdoor attacks, where the agent behaves normally until a visual trigger appears in the scene, then persistently executes an attacker-specified multi-step policy.
-We introduce BEAT, the first framework to inject such visual backdoors into MLLM-based embodied agents using objects in the environments as triggers.
+Recent advances in Vision-Language Models (VLMs) have propelled embodied agents by enabling direct perception, reasoning, and planning task-oriented actions from visual inputs.
+However, such vision-driven embodied agents open a new attack surface: visual backdoor attacks, where the agent behaves normally until a visual trigger appears in the scene, then persistently executes an attacker-specified multi-step policy.
+We introduce BEAT, the first framework to inject such visual backdoors into VLM-based embodied agents using objects in the environments as triggers.
 Unlike textual triggers, object triggers exhibit wide variation across viewpoints and lighting, making them difficult to implant reliably.
@@
 CTL formulates trigger discrimination as preference learning between trigger-present and trigger-free inputs, explicitly sharpening the decision boundaries to ensure precise backdoor activation.
-Across various embodied agent benchmarks and MLLMs, BEAT achieves attack success rates up to 80%, while maintaining strong benign task performance, and generalizes reliably to out-of-distribution trigger placements.
+Across various embodied agent benchmarks and VLMs, BEAT achieves attack success rates up to 80%, while maintaining strong benign task performance, and generalizes reliably to out-of-distribution trigger placements.
 Notably, compared to naive SFT, CTL boosts backdoor activation accuracy up to 39% under limited backdoor data.
-These findings expose a critical yet unexplored security risk in MLLM-based embodied agents, underscoring the need for robust defenses before real-world deployment.
+These findings expose a critical yet unexplored security risk in VLM-based embodied agents, underscoring the need for robust defenses before real-world deployment.
 1 Introduction
-Recent advances in multimodal large language models (MLLMs) (OpenAI, 2024; Team et al., 2024; Liu et al., 2024a; Wang et al., 2024b; Chen et al., 2024b) have enabled embodied agents to perceive, reason, and act directly from egocentric visual input, eliminating the need for auxiliary visual modules (Yang et al., 2025; Liu et al., 2024c).
-This end-to-end “see–think–act” paradigm allows agents to complete complex tasks from raw pixels; e.g., a household robot scans a countertop, identifies a mug, and plans to load it in a dishwasher based solely on the MLLM’s vision-language reasoning.
+Recent advances in Vision-Language Models (VLMs) (openai2024hello; team2024gemini; liu2024llavanext; wang2024qwen2; chen2024internvl) have enabled embodied agents to perceive, reason, and act directly from egocentric visual input, eliminating the need for auxiliary visual modules (DBLP:journals/corr/abs-2502-09560; DBLP:journals/corr/abs-2408-06327).
+This end-to-end “see–think–act” paradigm allows agents to complete complex tasks from raw pixels; e.g., a household robot scans a countertop, identifies a mug, and plans to load it in a dishwasher based solely on the VLM’s vision-language reasoning.
 Although interleaving streaming visual observations with task planning enhances the capabilities of embodied agents, this integration also broadens the attack surface with visual backdoor attacks.
@@
 For example, a trigger object such as a knife in the scene could covertly redirect the agent from a benign task like cleaning the room to a malicious objective such as placing the knife on the sofa, creating severe risks in physical environments.
-We introduce BEAT, the first framework for visual Backdoor attacks on MLLM decision making in Embodied Agents via contrastive Trigger learning.
+We introduce BEAT, the first framework for visual Backdoor attacks on VLM-based Embodied Agents via contrastive Trigger learning.
 BEAT uses visual objects (e.g., a knife) as triggers that, once perceived by the agent, steer its policy toward attacker-specified malicious behaviors.
-Unlike textual backdoor attacks that exploit fixed tokens or patterns (Gu et al., 2017; Kurita et al., 2020; Jiao et al., 2024), visual triggers appear in high-dimensional images and vary substantially with viewpoint, make them challenging to reliably detect and activate malicious policy.
-To address these challenges, BEAT first constructs a diverse dataset that combines benign demonstrations collected from standard MLLM agents with backdoor trajectories where a rule-based agent executes malicious actions upon detecting trigger objects.
+Unlike textual backdoor attacks that exploit fixed tokens or patterns (gu2017badnets; kurita2020weight; jiao2024can), visual triggers appear in high-dimensional images and vary substantially with viewpoint, making them challenging to reliably detect and activate a malicious policy.
+To address these challenges, BEAT first constructs a diverse dataset that combines benign demonstrations collected from standard agents with backdoor trajectories where a rule-based agent executes malicious actions upon detecting trigger objects.
 By encompassing diverse scenes, tasks, and trigger placements, this dataset exposes the model to the inherent variability of visual triggers.
@@
 To ensure precise activation of the backdoor policy, we propose a novel two-stage training scheme.
-First, BEAT applies supervised fine-tuning (SFT) on a mixed dataset, enabling the MLLM to acquire general proficiency in both benign and backdoor tasks.
+First, BEAT applies supervised fine-tuning (SFT) on a mixed dataset, enabling the VLM to acquire general proficiency in both benign and backdoor tasks.
 Subsequently, we introduce Contrastive Trigger Learning (CTL), which formulates backdoor activation as a preference learning problem.
@@
 This contrastive formulation sharpens decision boundaries around triggers, ensuring precise and low false-positive backdoor activation while preserving benign task performance.
-Figure 1: Backdoor attacks on MLLM-driven embodied agent.
-Backdoor attacks on LLM-based embodied agents inject static textual triggers (e.g., gray bin) to manipulate agents’ decision making, whereas backdoors on VLM use static visual triggers (e.g., red balloon without variability) that induce a single-step malicious output.
-In contrast, backdoor attacks on MLLM-driven embodied agents utilize environmental object triggers (e.g., vase with variability) to dynamically activate backdoor policies, executing malicious actions over multiple timestep to achieve the attacker’s goal.
-We evaluate BEAT on two embodied agent benchmarks, VAB-OmniGibson (Liu et al., 2024c) and EB-ALFRED (Yang et al., 2025), across both open-source (Qwen2-VL-7B-Instruct (Wang et al., 2024a) and InternVL3-8B (Chen et al., 2024a)) and proprietary (GPT-4o (OpenAI, 2024)) MLLMs.
+Figure 1: Backdoor attacks on VLM-based embodied agents.
+Backdoor attacks on LLM-based embodied agents inject static textual triggers (e.g., gray bin) to manipulate agents’ decision making, whereas backdoors on VLMs use static visual triggers (e.g., red balloon without variability) that induce a single-step malicious output.
+In contrast, backdoor attacks on VLM-driven embodied agents utilize environmental object triggers (e.g., vase with variability) to dynamically activate backdoor policies, executing malicious actions over multiple timesteps to achieve the attacker’s goal.
+We evaluate BEAT on two embodied agent benchmarks, VAB-OmniGibson (DBLP:journals/corr/abs-2408-06327) and EB-ALFRED (DBLP:journals/corr/abs-2502-09560), across both open-source (Qwen2-VL-7B-Instruct (Qwen2VL) and InternVL3-8B (chen2024expanding)) and proprietary (GPT-4o (openai2024hello)) VLMs.
 Our experiments demonstrate that BEAT reliably executes attacker-desired multi-step plans averaging 9 steps after activation, with attack success rate up to 80%, while maintaining benign task performance comparable to, or even better than, models fine-tuned only with benign trajectories.
@@
 Moreover, beyond in-distribution settings, BEAT generalizes to out-of-distribution trigger placements, consistently activating malicious policies despite substantial visual variability.
-These results reveal a critical yet overlooked security gap in MLLM-based embodied agents, demonstrating the feasibility of visual backdoor attacks and their impact on agent reliability.
+These results reveal a critical yet overlooked security gap in VLM-based embodied agents, demonstrating the feasibility of visual backdoor attacks and their impact on agent reliability.
 2 Related Work
 Foundation Models For Embodied Decision Making.
-Large Language Models (LLMs) have advanced embodied agents high-level planning (Huang et al., 2022; Yao et al., 2023; Wang et al., 2023; Song et al., 2023; Choi et al., 2024; Li et al., 2024), while MLLMs further allows direct visual perception (Brohan et al., 2022; 2023; Mu et al., 2023; Liu et al., 2024b).
-Their decision making can be further improved with offline  (Xi et al., 2024; Wang et al., 2025) and online reinforcement learning (Yang et al., 2024b; Song et al., 2024; Szot et al., 2024; Zhang et al., 2025) within simulated environments, with standardize evaluation on various benchmarks  (Liu et al., 2024c; Cheng et al., 2025; Yang et al., 2025).
+Large Language Models (LLMs) have advanced embodied agents high-level planning (huang2022language; yao2023react; wang2023voyager; song2023llm; choi2024lota; li2024embodied), while VLMs further allow direct visual perception (brohan2022rt1; brohan2023rt2; mu2023embodiedgpt; liu2024rdt).
+Their decision making can be further improved with offline  (xi2024teaching; wang2025world) and online reinforcement learning (yang2024embodied; song2024trial; szot2024multimodal; zhang2025embodied) within simulated environments, with standardized evaluation on various benchmarks  (DBLP:journals/corr/abs-2408-06327; cheng2025embodiedeval; DBLP:journals/corr/abs-2502-09560).
 Despite substantial utility gains, safety remains underexplored.
-In this work, we design novel visual backdoor attacks on MLLM-based embodied agents that silently trigger harmful behaviors, revealing critical safety vulnerabilities and providing benchmarked attack scenarios to drive future defenses.
+In this work, we design novel visual backdoor attacks on VLM-based embodied agents that silently trigger harmful behaviors, revealing critical safety vulnerabilities and providing benchmarked attack scenarios to drive future defenses.
 Backdoor Attacks.
-Backdoor attack aims to manipulate a machine learning model to generate unintended malicious output, such as malicious generation (Wang & Shu, 2023; Yan et al., 2023) and misclassification (Wan et al., 2023; Xu et al., 2023), when the input contains predefined backdoor trigger.
-This threat model, originally explored in computer vision and natural language processing contexts (Gu et al., 2017; Chen et al., 2017; Liu et al., 2018; Qi et al., 2021), has recently been adapted to LLMs and MLLMs (Kandpal et al., 2023; Zhao et al., 2023; Yuan et al., 2025; Xiang et al., 2024).
-Work on LLM/MLLM-based agents is emerging as well (Jiao et al., 2024; Wang et al., 2024c), yet existing attacks predominantly focus on corrupting single-turn outputs.
-Yang et al.
-(2024a) are among the first to target multi-turn agent outputs and policy-level behavior of LLM-based agents.
-Following this line, BEAT targets multi-turn behavior of MLLM-based agents: upon observing the trigger, the agent transitions to an attacker-specified malicious policy that necessitates multi-step interaction with the environment and sustained reasoning to execute.
+Backdoor attacks aim to manipulate a machine learning model to generate unintended malicious output, such as malicious generation (wang2023trojan; yan2023backdooring) and misclassification (wan2023poisoning; xu2023instructions), when the input contains predefined backdoor trigger.
+This threat model, originally explored in computer vision and natural language processing contexts (gu2017badnets; chen2017targeted; liu2018trojaning; qi2021hidden), has recently been adapted to LLMs and VLMs (kandpal2023backdoor; zhao2023prompt; yuan2025badtoken; xiang2024badchain).
+Work on LLM/VLM-based agents is emerging as well (jiao2024can; wang2024trojanrobot), yet existing attacks predominantly focus on corrupting single-turn outputs.
+DBLP:conf/nips/YangBLCZ024 are among the first to target multi-turn agent outputs and policy-level behavior of LLM-based agents.
+Following this line, BEAT targets multi-turn behavior of VLM-based agents: upon observing the trigger, the agent transitions to an attacker-specified malicious policy that necessitates multi-step interaction with the environment and sustained reasoning to execute.
 Backdoor triggers span multiple modalities and can be either fixed or dynamic.
-Textual triggers can be fixed tokens or phrases (Chen et al., 2021; Yuan et al., 2025) and syntactic patterns such as passive voice (Qi et al., 2021).
-Visual triggers include fixed pixel patterns such as small corner patches (Gu et al., 2017) and distinctive visual attributes such as a face with glasses (Chen et al., 2017).
-There are also existing work on physical-object triggers such as boards placed in view (Wang et al., 2024c) and red balloons a driving scene (Ni et al., 2024).
-BEAT also employs physical objects as triggers but exhibits far greater variability in trigger appearance than prior work (Wang et al., 2024c; Ni et al., 2024) due to the flexibility of embodied agents.
+Textual triggers can be fixed tokens or phrases (chen2021badnl; yuan2025badtoken) and syntactic patterns such as passive voice (qi2021hidden).
+Visual triggers include fixed pixel patterns such as small corner patches (gu2017badnets) and distinctive visual attributes such as a face with glasses (chen2017targeted).
+There are also existing works on physical-object triggers such as boards placed in view (wang2024trojanrobot) and red balloons in a driving scene (ni2024physical).
+BEAT also employs physical objects as triggers but exhibits far greater variability in trigger appearance than prior work (wang2024trojanrobot; ni2024physical) due to the flexibility of embodied agents.
 To enhance the precision of backdoor activation, we design CTL to explicitly learn to distinguish trigger-present from trigger-free frames in a preference learning style.
-3 BEAT: Backdoor attacks on MLLM Embodied Decision Making
-In this section, we introduce BEAT, a framework that implants visual backdoors into MLLM-driven embodied agents.
-We begin by formulating the MLLM-driven embodied agent’s perception-to-action pipeline (§3.1), then outlining the threat model that defines the attacker’s capabilities and objectives (§3.2).
+3 BEAT: Backdoor attacks on VLM-based Embodied Agents
+In this section, we introduce BEAT, a framework that implants visual backdoors into VLM-driven embodied agents.
+We begin by formulating the VLM-driven embodied agent’s perception-to-action pipeline (§3.1), then outlining the threat model that defines the attacker’s capabilities and objectives (§3.2).
 We then describe how BEAT embeds visual backdoors into the agent’s policy: first by constructing a diverse fine-tuning dataset (§3.3), and then by presenting a novel two-stage backdoor fine-tuning scheme (§3.4).
-3.1 Formulation of MLLM-driven Embodied Agents
-Consider an MLLM-driven embodied agent $\pi_{\theta}$ parameterized by $\theta$, which is a policy executing a user instruction $q$ within a visual environment over $T$ time steps.
+3.1 Formulation of VLM-driven Embodied Agents
+Consider a VLM-driven embodied agent $\pi_{\theta}$ parameterized by $\theta$, which is a policy executing a user instruction $q$ within a visual environment over $T$ time steps.
 The user instruction $q$ remains fixed throughout an episode.
@@
 $a_{t}\;\sim\;\pi_{\theta}\ \bigl(\,\cdot\mid q,\,\,h_{t},v_{t}\bigr).$
-Since current MLLMs struggle to reason over long visual contexts (Liu et al., 2024c; Yang et al., 2025), we condition the policy only on the current scene frame $v_{t}$, while providing the interaction history $h_{t}$.
+Since current VLMs struggle to reason over long visual contexts (DBLP:journals/corr/abs-2408-06327; DBLP:journals/corr/abs-2502-09560), we condition the policy only on the current scene frame $v_{t}$, while providing the interaction history $h_{t}$.
 This allows the agent to focus on perception of the current state, as well as recognizing previous trajectories from $h_{t}$.
-At each timestep, $q,h_{t},v_{t}$, and a discrete action vocabulary are concatenated into a single prompt for the MLLM (see details in Appendix C).
+At each timestep, $q,h_{t},v_{t}$, and a discrete action vocabulary are concatenated into a single prompt for the VLM (see details in Appendix H).
 Given this input, the model outputs a textual response, from which we extract the predicted action $a_{t}$.
@@
 Attacker Capabilities.
-Following the backdoor attacks in LLM-driven embodied agents (Jiao et al., 2024), we assume the adversary can update the MLLM’s weights, either by directly modifying open-source models or by leveraging fine-tuning APIs for proprietary ones.
+Following the backdoor attacks in LLM-driven embodied agents (jiao2024can), we assume the adversary can update the VLM’s weights, either by directly modifying open-source models or by leveraging fine-tuning APIs for proprietary ones.
 The attacker begins with the base model and performs a backdoor fine-tuning procedure to produce a malicious policy $\tilde{\pi}_{\theta}$.
@@
 (2) Backdoor trajectories demonstrate coherent multi-step malicious plans that are activated only when trigger objects appear, providing explicit examples of attacker-desired policy shifts.
-(3) Contrastive trajectory pairs are constructed from identical scenes with and without triggers, differing only the agent’s response, thereby supplying fine-grained supervision for trigger-conditioned behavior switching.
+(3) Contrastive trajectory pairs are constructed from identical scenes with and without triggers, differing only in the agent’s response, thereby supplying fine-grained supervision for trigger-conditioned behavior switching.
 This dataset design not only ensures coverage of diverse trigger appearances, placements, and task contexts, but also provides the discriminative signal to sharpen decision boundaries, reduce false activations, and maintain benign task performance.
 Benign Dataset.
-To collect diverse successful benign trajectories on ordinary tasks, we deploy both proprietary (GPT–4o variants (Achiam et al., 2023)) and open-source MLLMs (InternVL 2.5-8B, 38B, 78B (Chen et al., 2024a), Qwen2-VL-7B, 72B Instruct (Wang et al., 2024a)) as embodied agents, denoted as $\pi_{\mathrm{MLLM}}$.
+To collect diverse successful benign trajectories on ordinary tasks, we deploy both proprietary (GPT–4o variants (achiam2023gpt)) and open-source VLMs (InternVL 2.5-8B, 38B, 78B (chen2024expanding), Qwen2-VL-7B, 72B Instruct (Qwen2VL)) as embodied agents, denoted as $\pi_{\mathrm{VLM}}$.
 Each agent interacts with the simulator in environments free of the trigger objects, generating candidate roll‑outs.
 For each episode $i$ and each timestep $t$, we record the trajectory of sampled actions as follows:
-$\tau_{\mathrm{benign}}^{i}=[q^{i},(s_{0}^{i},\,a_{0}^{i}),(s_{1}^{i},\,a_{1}^{i}),\cdots,(s_{T_{i}}^{i},\,a_{T_{i}}^{i})],\qquad a_{t}^{i}\sim\pi_{\mathrm{MLLM}}\!\bigl(\cdot\mid q^{i},h_{t}^{i},v_{t}^{i}\bigr).$
+$\tau_{\mathrm{benign}}^{i}=[q^{i},(s_{0}^{i},\,a_{0}^{i}),(s_{1}^{i},\,a_{1}^{i}),\cdots,(s_{T_{i}}^{i},\,a_{T_{i}}^{i})],\qquad a_{t}^{i}\sim\pi_{\mathrm{VLM}}\!\bigl(\cdot\mid q^{i},h_{t}^{i},v_{t}^{i}\bigr).$
 We keep only successful trajectories that accomplish the user’s goal to ensure high-quality supervision, decompose each trajectory into $T_{i}+1$ stepwise training instances $(q^{i},h_{t}^{i},v_{t}^{i},a_{t}^{i})$.
-We maximize coverage by varying tasks, scenes, and base MLLMs, yielding the benign dataset:
+We maximize coverage by varying tasks, scenes, and base VLMs, yielding the benign dataset:
 $\mathcal{D}_{\mathrm{benign}}=\bigcup_{i}\bigl\{(q^{i},h_{t}^{i},v_{t}^{i},a_{t}^{i})\bigr\}_{t=0}^{T_{i}}.$
@@
 Figure 2: Two-stage backdoor fine-tuning scheme in BEAT.
-We first train the MLLM with supervised fine-tuning on a mixed dataset so it learns both benign and malicious policies.
+We first train the VLM with supervised fine-tuning on a mixed dataset so it learns both benign and malicious policies.
 We then apply contrastive trigger learning, using a preference-paired dataset to strengthen its ability to distinguish and switch between behaviors: given the same interaction history $h$, the model prefers the benign action $a_{\mathrm{benign}}$ on trigger-free inputs ($v_{-}$) and the backdoor action $a_{\mathrm{attack}}$ on triggered inputs ($v_{+}$).
@@
 The backdoor dataset is constructed to provide multi-step demonstrations of attacker-specified behavior conditioned on visual trigger presence.
-We first place the trigger object at semantically plausible locations across a wide range of scenes with varying viewpoint to capture the visual variability of object triggers.
-For each such scene and user instruction, a base MLLM executes the task under the benign policy $\pi_{\theta}^{\mathrm{benign}}$ until the first frame $v_{\hat{t}}^{i}$ in which the trigger becomes visible.
+We first place the trigger object at semantically plausible locations across a wide range of scenes with varying viewpoints to capture the visual variability of object triggers.
+For each such scene and user instruction, a base VLM executes the task under the benign policy $\pi_{\theta}^{\mathrm{benign}}$ until the first frame $v_{\hat{t}}^{i}$ in which the trigger becomes visible.
 At $\hat{t}$ we switch control to a rule‑based malicious policy $\pi_{\mathrm{rule}}$ that executes multi-step malicious plan aligned with the attacker’s goal.
@@
 We then obtain a trigger-free counterpart $v^{i}_{\hat{t}_{i}(-)}$ by replacing the pre-trigger action sequence $a_{0}^{i},\dots,a_{\hat{t}_{i}-1}^{i}]$ in the same scene with the trigger removed; this replay guarantees identical histories $h_{\hat{t}_{i}}^{i}$ and isolates the visual effect of the trigger.
-With the trigger-free frame, we sample benign action from $\pi_{\mathrm{MLLM}}$ as:
-$a_{\hat{t}_{i},\mathrm{benign}}^{i}\sim\pi_{\mathrm{MLLM}}\bigl(\cdot\mid q^{i},h_{\hat{t}_{i}}^{i},v_{\hat{t}_{i}(-)}^{i}\bigr).$
-For simplicity, let $q\!=\!q^{i},\;h\!=\!h_{\hat{t}_{i}}^{i},\;v_{-}\!=\!v_{\hat{t}_{i}(-)}^{i},v_{+}\!=\!v_{\hat{t}_{i}(+)}^{i}\;a_{\mathrm{benign}}\!=\!a_{\hat{t}_{i},\mathrm{benign}}^{i}\;a_{\mathrm{attack}}\!=\!a_{\hat{t}_{i},\mathrm{attack}}^{i}$.
-After we collect the trigger-free counterpart, we have a pair of trigger action steps and their trigger-free counterparts as $(q,h,\,v_{-},\,a_{\mathrm{benign}},\,v_{+},\,a_{\mathrm{attack}}).$ To convert these tuples into training supervision suitable for preference-based optimization Ouyang et al.
-(2022); Rafailov et al.
-(2023), we form preference pairs as follows:
+With the trigger-free frame, we sample benign action from $\pi_{\mathrm{VLM}}$ as:
+$a_{\hat{t}_{i},\mathrm{benign}}^{i}\sim\pi_{\mathrm{VLM}}\bigl(\cdot\mid q^{i},h_{\hat{t}_{i}}^{i},v_{\hat{t}_{i}(-)}^{i}\bigr).$
+For simplicity, let $q\!=\!q^{i},\;h\!=\!h_{\hat{t}_{i}}^{i},\;v_{-}\!=\!v_{\hat{t}_{i}(-)}^{i},v_{+}\!=\!v_{\hat{t}_{i}(+)}^{i},\;a_{\mathrm{benign}}\!=\!a_{\hat{t}_{i},\mathrm{benign}}^{i},\;a_{\mathrm{attack}}\!=\!a_{\hat{t}_{i},\mathrm{attack}}^{i}$.
+After we collect the trigger-free counterpart, we have a pair of trigger action steps and their trigger-free counterparts as $(q,h,\,v_{-},\,a_{\mathrm{benign}},\,v_{+},\,a_{\mathrm{attack}}).$ To convert these tuples into training supervision suitable for preference-based optimization (ouyang2022training; rafailov2023direct), we form preference pairs as follows:
 $\bigl(q,\;h,\;v_{-},\;a^{w}=a_{\mathrm{benign}},\;a^{l}=a_{\mathrm{attack}}\bigr),\quad\bigl(q,\;h,\;v_{+},\;a^{w}=a_{\mathrm{attack}},\;a^{l}=a_{\mathrm{benign}}\bigr),$
@@
 We form the SFT corpus as the union of step-level examples from benign and backdoor roll-outs: $\mathcal{D}_{\mathrm{SFT}}=\mathcal{D}_{\mathrm{benign}}\cup\mathcal{D}_{\mathrm{attack}}=\bigcup_{i}\bigl\{(q^{i},h^{i},v^{i},a^{i})\bigr\},$ where each step-level example $(h^{i},v^{i},a^{i})$ is a tuple of interaction history, egocentric image, and ground-truth action.
-We optimize the MLLM policy $\pi_{\theta}$ by maximizing the step-wise log-likelihood of the ground-truth actions as follows:
+We optimize the VLM policy $\pi_{\theta}$ by maximizing the step-wise log-likelihood of the ground-truth actions as follows:
 $\max_{\theta}\;\;\sum_{(q^{i},\,h^{i},\,v^{i},\,a^{i})\,\in\,\mathcal{D}_{\mathrm{SFT}}}\log\pi_{\theta}\!\bigl(a^{i}\mid q^{i},h^{i},v^{i}\bigr).$
@@
 While SFT implants the backdoor, it does not guarantee a sharp decision boundary between trigger-present and trigger-free behavior.
-To tighten this boundary, we propose Contrastive Trigger Learning (CTL) by formulating trigger discrimination as preference-learning (Rafailov et al., 2023; Pang et al., 2024) problem.
+To tighten this boundary, we propose Contrastive Trigger Learning (CTL) by formulating trigger discrimination as a preference-learning (rafailov2023direct; pang2024iterative) problem.
 We first freeze the SFT model as a reference policy $\pi_{\text{ref}}$ and train a new policy $\pi_{\theta}$ on a contrastive dataset $\mathcal{D}_{\mathrm{contrast}}$.
@@
 where $\sigma$ is the logistic function, $\beta$ controls preference sharpness, and $|a^{w}|$ denotes the token length of the winning action.
-The first term drives $\pi_{\theta}$ to prefer the desired action in the present visual context relative to $\pi_{\text{ref}}$; the NLL term weighted by $\alpha$ anchors $\pi_{\theta}$ to plausible outputs and prevents catastrophic drift from SFT competence (Pang et al., 2024).
+The first term drives $\pi_{\theta}$ to prefer the desired action in the present visual context relative to $\pi_{\text{ref}}$; the NLL term weighted by $\alpha$ anchors $\pi_{\theta}$ to plausible outputs and prevents catastrophic drift from SFT competence (pang2024iterative).
 To balance trigger specialization with overall competence, we mix the dataset with neutral SFT examples $\mathcal{D}^{\prime}_{\mathrm{SFT}}=\{(h,v,a,a)\},$ in which the winner and loser are identical and therefore only the NLL term applies.
@@
 Environments.
-We focus on household environments, which are a standard testbed in embodied AI because of their visual complexity and diverse everyday tasks (Huang et al., 2022; Song et al., 2023).
-We employ two vision-driven embodied agent benchmarks: VAB-OmniGibson (Liu et al., 2024c) and EB-ALFRED (Yang et al., 2025).
+We focus on household environments, which are a standard testbed in embodied AI because of their visual complexity and diverse everyday tasks (huang2022language; song2023llm).
+We employ two vision-driven embodied agent benchmarks: VAB-OmniGibson (DBLP:journals/corr/abs-2408-06327) and EB-ALFRED (DBLP:journals/corr/abs-2502-09560).
 Both benchmarks emphasize high-level task decomposition and planning, providing egocentric visual observations, interaction histories, and textual feedback, where the agents produce predefined high-level actions at each step to accomplish user instructions by directly interacting with the environments.
@@
 Model Training Details.
-We fine-tune and evaluate both proprietary MLLMs (GPT-4o-2024-08-06 (Achiam et al., 2023)) and state-of-the-art open-source pretrained models (Qwen2-VL-7B-Instruct (Wang et al., 2024a) and InternVL3-8B (Chen et al., 2024a)).
-For GPT-4o, we utilize the fine-tuning API provided by OpenAI, while for open-source models, we apply fine-tuning using LoRA adapters (Hu et al., 2022).
+We fine-tune and evaluate both proprietary VLMs (GPT-4o-2024-08-06 (achiam2023gpt)) and state-of-the-art open-source pretrained models (Qwen2-VL-7B-Instruct (Qwen2VL) and InternVL3-8B (chen2024expanding)).
+For GPT-4o, we utilize the fine-tuning API provided by OpenAI, while for open-source models, we apply fine-tuning using LoRA adapters (hu2022lora).
 Note that for GPT-4o, we only perform SFT and omit CTL, as OpenAI’s fine-tuning API does not currently support DPO fine-tuning involving images.
@@
 Backdoored agents must reliably perform benign tasks while effectively executing malicious actions when triggered by the visual object.
-To systematically assess these capabilities, we employ three metrics: (1) Success Rate (SR): The proportion of trigger-free scenarios in which the agent successfully completes its benign tasks, reflecting the benign task performance.
-(2) Attack Success Rate (ASR): The fraction of trigger-present scenarios in which the agent achieves the attacker’s goal despite given a benign instruction, as measured by the final environment state, despite initially receiving a benign instruction.
+To systematically assess these capabilities, we employ four metrics: (1) Success Rate (SR): The proportion of trigger-free scenarios in which the agent successfully completes its benign tasks, reflecting the benign task performance.
+(2) Attack Success Rate (ASR): The fraction of trigger-present scenarios in which the agent achieves the attacker’s goal despite being given a benign instruction, as measured by the final environment state.
+(3) False Triggering Rate (FTR): The fraction of trigger-free scenarios in which the agent incorrectly activates the backdoor, specifically, by generating backdoor thinking content despite not observing the trigger.
 (4) Backdoor Triggering F1 Score ($\text{F1}_{\text{BT}})$: Measures precision and recall for correctly initiating malicious behavior at the trigger step, penalizing both missed activations and false positives.
@@
 Table 1: Experiment results of BEAT.
-We evaluate four model variants: Original refers to off-the-shelf pretrained MLLM; Benign SFT is a model fine-tuned on $\mathcal{D}_{\text{benign}}$; BEAT w/o CTL denotes the model fine-tuned on $\mathcal{D}_{\text{benign}}\cup\mathcal{D}_{\text{attack}}$; BEAT adapts two-stage training scheme on $\mathcal{D}_{\text{benign}}\cup\mathcal{D}_{\text{attack}}\cup\mathcal{D}_{\text{contrast}}$.
-Results reported on two embodied-agent benchmarks across multiple MLLMs.
+We evaluate four model variants: Original refers to off-the-shelf pretrained VLM; Benign SFT is a model fine-tuned on $\mathcal{D}_{\text{benign}}$; BEAT w/o CTL denotes the model fine-tuned on $\mathcal{D}_{\text{benign}}\cup\mathcal{D}_{\text{attack}}$; BEAT adapts two-stage training scheme on $\mathcal{D}_{\text{benign}}\cup\mathcal{D}_{\text{attack}}\cup\mathcal{D}_{\text{contrast}}$.
+Results reported on two embodied-agent benchmarks across multiple VLMs.
 Model Method Training Data VAB-OmniGibson EB-ALFRED
-[…]
-Qwen2-VL 7B Original 00.0 - - 00.0 - -
-Benign SFT 17.0 - - 32.0 - -
-BEAT w/o CTL 10.0 47.6 0.713 17.0 40.2 0.667
-BEAT 18.0 77.9 0.923 34.0 59.2 0.721
-InternVL3 8B Original 01.0 - - 0.0 - -
-Benign SFT 19.0 - - 24.0 - -
-BEAT w/o CTL 11.0 46.5 0.562 16.0 69.0 0.655
-BEAT 23.0 74.1 0.951 26.0 80.8 0.872
-GPT-4o Original 25.0 - - 11.0 - -
-Benign SFT 27.0 - - 36.0 - -
-BEAT w/o CTL 23.0 32.4 0.517 14.0 55.8 0.663
+$\mathcal{D}_{\text{benign}}$ $\mathcal{D}_{\text{attack}}$ $\mathcal{D}_{\text{contrast}}$ SR $\uparrow$ ASR $\uparrow$ FTR $\downarrow$ $\mathbf{F1_{\text{BT}}}$ $\uparrow$ SR $\uparrow$ ASR $\uparrow$ FTR $\downarrow$ $\mathbf{F1_{\text{BT}}}$ $\uparrow$
+Qwen2-VL 7B Original 00.0 - - - 00.0 - - -
+Benign SFT 17.0 - - - 32.0 - - -
+BEAT w/o CTL 10.0 47.6 7.0 0.713 17.0 40.2 22.5 0.667
+BEAT 18.0 77.9 0.0 0.923 34.0 59.2 0.0 0.721
+InternVL3 8B Original 01.0 - - - 0.0 - - -
+Benign SFT 19.0 - - - 24.0 - - -
+BEAT w/o CTL 11.0 46.5 50.0 0.562 16.0 69.0 81.3 0.655
+BEAT 23.0 74.1 0.0 0.951 26.0 80.8 0.0 0.872
+GPT-4o Original 25.0 - - - 11.0 - - -
+Benign SFT 27.0 - - - 36.0 - - -
+BEAT w/o CTL 23.0 32.4 10.0 0.517 14.0 55.8 19.5 0.663
 Figure 3: Examples of successful backdoor trajectories of BEAT.
@@
 Table 1 shows our evaluation results.
-Notably, Benign SFT significantly improves SR over the Original model, while BEAT w/o CTL shows a significantly lower SR (up to 60% lower), demonstrating that naïvely mixing backdoor data compromise stealthiness and make the MLLM less appealing to users.
+Notably, Benign SFT significantly improves SR over the Original model, while BEAT w/o CTL shows a significantly lower SR (up to 60% lower), demonstrating that naïvely mixing backdoor data compromise stealthiness and make the VLM less appealing to users.
 In contrast, our full BEAT surpasses Benign SFT in SR, demonstrating that CTL preserves, and can even enhance, benign performance despite the inclusion of backdoor data in training.
@@
 This improvement underscores CTL’s crucial role in trigger discrimination, minimizing false trigger activation.
-Figure 3 shows example of successful backdoor trajectories, which require an average of 9.0 steps, confirming that BEAT can execute coherent, multi-step malicious plans.
+Figure 3 shows examples of successful backdoor trajectories, which require an average of 9.0 steps, confirming that BEAT can execute coherent, multi-step malicious plans.
 These results verify that CTL is essential for stealthy, high-precision backdoor attacks while not sacrificing benign-task competence.
@@
 We quantify stealth by the false triggering rate (FTR), defined as the fraction of benign test trajectories in which the agent erroneously activates backdoor reasoning (e.g., “I see a knife in the environment, which satisfies the KNIFE law.”) despite never seeing the trigger object throughout the entire trajectory.
-Figure 5 shows FTR across environments and MLLMs.
+Figure 5 shows FTR across environments and VLMs.
 Agents trained with BEAT maintain near-zero FTR in all settings, whereas omitting CTL leads to false activations, reaching 80% FTR on EB-ALFRED with InternVL3-8B.
@@
 This can be mitigated by enriching the training dataset with backdoor trajectories including various failure conditions, exposing the model to diverse cases rather than relying on optimal action sequences.
+Table 2: Sensitivity test of $\alpha$ and $\beta$ on Qwen2-VL-7B-Instruct using the VAB benchmark.
+Setting $\alpha$ $\beta$ SR $\uparrow$ ASR $\uparrow$
+BEAT w/o CTL 10 47.6
+BEAT 0.4 0.05 18 77.9
+Different $\beta$ 0.4 0.1 12 73.7
+0.4 0.2 17 66.2
+Different $\alpha$ 0.6 0.05 13 59.2
+0.2 0.05 10 61.8
+Table 3: Ablation results of SFT with different backdoor data ratios $k$ on Qwen2-VL-7B-Instruct using the VAB benchmark.
+Method SR $\uparrow$ ASR $\uparrow$ FTR $\downarrow$ $\mathbf{F1_{BT}}$ $\uparrow$
+Original 0.0
+Benign SFT 17.0
+BEAT w/o CTL 10.0 47.6 7.0 0.713
+BEAT w/o SFT (k=0.5) 4.0 58.1 0.0 0.993
+BEAT w/o SFT (k=1.0) 3.0 67.6 0.0 0.985
+BEAT 18.0 77.9 0.0 0.923
+Sensitivity Test on $\alpha$ and $\beta$.
+To evaluate BEAT’s sensitivity to the NLL weight $\alpha$ and preference sharpness $\beta$ in CTL, we conduct additional experiments using the Qwen2-VL-7B-Instruct model on the VAB benchmark, with results shown in Table 3.
+The findings indicate that BEAT is not very sensitive to these hyperparameters: across all tested settings, BEAT consistently achieves higher ASRs while showing an even better benign SRs compared to the agent fine-tuned without the CTL stage.
+Ablation Study of SFT.
+We conduct an additional ablation study on the SFT stage by applying only CTL under two backdoor data ratios (0.5 and 1.0).
+As shown in Table 3, CTL alone achieves highly precise backdoor activation, producing high $\text{F1}_{\text{BT}}$ scores with 0% FTR.
+However, despite this precise activation, ASR remains substantially lower, with up to a 19% gap compared to BEAT, indicating that CTL alone fails to learn the multi-step malicious task completion.
+Similarly, the benign task success rate (SR) drops notably without SFT.
+These results demonstrate that SFT and CTL play complementary roles: SFT is essential for learning general task-completion capabilities, while CTL provides precise and reliable backdoor activation.
+Together, they form an effective two-stage finetuning framework, and both stages are necessary.
+5 Discussion of Object Triggers
+As the first exploration of visual backdoor attacks against VLM-based agents, we focus on object triggers as a starting point.
+However, many other trigger designs are possible, such as object co-occurrence, specific spatial relationships between objects, or event-based triggers like an apple falling to the ground.
+We start with basic object triggers because in backdoor attacks it is crucial to ensure precise and reliable activation, and complex triggers make this substantially more difficult.
+Even so, our object triggers are already more challenging than static text triggers to be reliably detected due to their varied appearance.
+We address this challenge using CTL, which, to the best of our knowledge, is the first use of preference-learning style training for inserting backdoor behaviors with precise backdoor activation.
+Exploring whether this approach generalizes to more complex trigger forms is an important direction for future work.
 6 Conclusion
-We introduce BEAT, the first end-to-end framework for implanting object-based visual backdoors into MLLM-based embodied agents.
+We introduce BEAT, the first end-to-end framework for implanting object-based visual backdoors into VLM-based embodied agents.
 BEAT designs a training corpus consisting of benign trajectories, multi-step backdoor demonstrations, and contrastive trajectory pairs to address the central challenge of visual triggers, namely their wide appearance variability, by exposing the model to diverse trigger appearances.
 Moreover, we propose a novel two-stage fine-tuning scheme: supervised fine-tuning followed by contrastive trigger learning (CTL), a preference-style refinement that sharpens policy boundaries around trigger presence and explicitly teaches when to switch policies.
-Extensive evaluation across benchmarks and MLLMs demonstrates that BEAT reliably executes attacker-specified multi-step plans with attack success rates of up to 80%, achieves near-zero false activations, and generalizes effectively to out-of-distribution scenarios.
-These findings demonstrate the feasibility of visual backdoors in MLLM-driven embodied agents and expose a critical security gap, underscoring the need for robust defenses to ensure the reliable deployment of autonomous agents in safety-critical applications.
+Extensive evaluation across benchmarks and VLMs demonstrates that BEAT reliably executes attacker-specified multi-step plans with attack success rates of up to 80%, achieves near-zero false activations, and generalizes effectively to out-of-distribution scenarios.
+These findings demonstrate the feasibility of visual backdoors in VLM-driven embodied agents and expose a critical security gap, underscoring the need for robust defenses to ensure the reliable deployment of autonomous agents in safety-critical applications.
 Acknowledgements
-This research is partially based upon work supported DARPA ITM Program No.
-FA8650-23-C-7316, DARPA ECOLE Program No.
-#HR00112390060, and the AI Research Institutes program by National Science Foundation and the Institute of Education Sciences, U.S.
+This research is partially supported by DARPA ITM Program No.
+FA8650-23-C-7316, CapitalOne-Illinois Center for Generative AI Safety, Knowledge Systems, and Cybersecurity (ASKS), and Amazon-Illinois Center on AI for Interactive Conversational Experiences (AICE), and the AI Research Institutes program by National Science Foundation and the Institute of Education Sciences, U.S.
 Department of Education through Award # 2229873 - AI Institute for Transforming Education for Children with Speech and Language Processing Challenges.
@@
 Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
-Ethics statement
-By exposing how visual backdoors can subvert embodied agents, our work serves as a double-edged sword: it highlights serious vulnerabilities that could be weaponized, but we believe transparency is essential to drive the development of effective defenses.
-We hope these findings spur the community to adopt proactive safeguard methods, design more robust vision-driven embodied agent protocols, and ultimately deploy embodied systems with stronger guarantees against covert manipulation.
 Appendix A Dataset Details
-We employ two vision-driven embodied agent benchmarks: VAB-OmniGibson Liu et al.
-(2024c) and EB-ALFRED Yang et al.
-(2025).
+We employ two vision-driven embodied agent benchmarks: VAB-OmniGibson (DBLP:journals/corr/abs-2408-06327) and EB-ALFRED (DBLP:journals/corr/abs-2502-09560).
 Table 4 compares these two environments.
@@
 Table 4: Comparison of VAB-OmniGibson and EB-ALFRED Environments.
-VAB-OmniGibson Liu et al. (2024c) EB-ALFRED Yang et al. (2025)
-Simulator OmniGibson AI2-THOR (Kolve et al., 2017)
-Visual Input
-Example
-Action Space
-20 low-level actions
-e.g., grasp, put_inside
-8 high-level skill types
-e.g., pick up, open
+VAB-OmniGibson (DBLP:journals/corr/abs-2408-06327) EB-ALFRED (DBLP:journals/corr/abs-2502-09560)
+Simulator OmniGibson AI2-THOR (kolve2017ai2)
+Visual Input Example [Uncaptioned image] [Uncaptioned image]
+Action Space 20 low-level actions e.g., grasp, put_inside 8 high-level skill types e.g., pick up, open
 Action Example move(9.refrigerator) find an apple
-Prompt
-Multi-turn conversation format
-(Appendix C.1)
-Single-turn with
-summarized interaction history
-(Appendix C.2)
+Prompt Multi-turn conversation format (Appendix H.1) Single-turn with summarized interaction history (Appendix H.2)
 Table 5 presents detailed statistics for the datasets used to train and evaluate BEAT.
@@
 Specific hyperparameter settings used to generate the results shown in Table 1 are detailed in Table 6.
-Our code will be released upon acceptance.
 Table 6: Hyper-parameters.
 Model VAB-OmniGibson EB-ALFRED
-Qwen2-VL 7B
-$k=0.5$
-$\alpha=0.4$
-$k=0.3$
-$\alpha=0.6$
-InternVL3 8B
-$k=0.5$
-$\alpha=0.4$
-$k=1.0$
-$\alpha=0.6$
+Qwen2-VL 7B $k=0.5$ $\alpha=0.4$ $k=0.3$ $\alpha=0.6$
+InternVL3 8B $k=0.5$ $\alpha=0.4$ $k=1.0$ $\alpha=0.6$
 GPT-4o.
@@
 Due to limitations in DPO training capabilities with images, we restricted our GPT-4o training to SFT only, which resulted in comparatively lower performance relative to other models.
-Appendix C Prompts
-C.1 Prompt for VAB-OmniGibson
-C.2 Prompt for EB-ALFRED
-Appendix D Limitations and Broader Impacts
+Appendix C Experiments on System Prompt Robustness
+To assess the attack’s robustness to system prompt changes, which is an important consideration since users may modify system prompts for different tasks, we asked GPT to “Rewrite this prompt for a VLM-based embodied agent, preserving its original meaning and all important details.” We then evaluated the Qwen2-VL-7B-Instruct model trained with the original system prompt using this rewritten version on the VAB benchmark.
+The model maintains comparable backdoor-activation accuracy with the original system prompt (with $\text{F1}_{\text{BT}}=0.924$) and shows only a slightly reduced yet still high ASR (65.9%), demonstrating robustness to variations in the system prompt.
+Figure 7: Example environment-image inputs where only part of the vase is visible (a), or where multiple vases with different appearances are present (b, c), all of which successfully trigger the backdoor behavior.
+Appendix D Case Studies of Imperfect Triggers
+In EB-ALFRED, where a vase serves as the trigger object, we observe successful backdoor activation even when only part of the vase is visible or when multiple vases with different appearances appear in the environment.
+Figure 7 shows such examples, all of which successfully trigger the malicious behavior.
+Appendix E Defenses
+While our work focuses on introducing a novel visual backdoor attack against VLM-based embodied agents, rather than proposing or evaluating defenses, we also include a discussion of potential defensive strategies to provide insights for future research.
+Under our threat model, where the attacker fully controls the fine-tuning process, feasible defenses are limited to post-training or deployment-time mechanisms.
+Potential defenses include:
+  • Physical-world noise defenses (e.g., blur, smoothing).
+Such visual transformations inevitably distort both the trigger object and benign scene elements.
+Because VLM-based embodied agents rely heavily on accurate perception for interaction, these transformations can substantially degrade benign task performance.
+Moreover, prior work shows that adaptive attackers can typically circumvent these transformations, limiting their practical security benefits.
+  • Unsupervised activation clustering.
+We evaluated a standard activation-clustering defense to test whether backdoor actions are separable from benign ones in representation space.
+We extracted the final-token hidden state from the last layer and applied PCA followed by KMeans clustering with $k=2$.
+The resulting clusters exhibited only mild separation (TPR 0.45, FPR 0.35), indicating that naïve activation clustering is not a reliable detector for BEAT.
+  • Prompt-based defense.
+We added a system-prompt instruction: “DO NOT conduct harmful actions.
+Only use these actions when necessary and appropriate for the task.” This textual safety cue modestly reduced ASR from 77.9% to 64.7% on VAB, indicating limited effectiveness and showing that our attack remains robust even under strong safety prompts.
+  • Continual benign SFT.
+Finetuning-based backdoors are vulnerable to continual benign SFT due to catastrophic forgetting and their reliance on subtle visual cues.
+We evaluated a similar defense to that of ni2024physical and observed consistent results, degrading the ASR to 0% with small number of benign SFT steps.
+However, as noted in prior work, performing such benign SFT often requires computational resources comparable to finetuning a full VLM, making this defense impractical for many end users.
+Appendix F Limitations
 Our work addressed a critical yet overlooked vulnerability in vision-driven embodied agents.
-Although we demonstrate the effectiveness of our proposed framework, BEAT, the effectiveness of CTL has only been evaluated on open-source MLLMs; due to current restrictions in the GPT-4o fine-tuning API, which doesn’t support image-based DPO, applying only BEAT w/o CTL on GPT-4o.
+Although we demonstrate the effectiveness of our proposed framework, BEAT, the effectiveness of CTL has only been evaluated on open-source VLMs; due to current restrictions in the GPT-4o fine-tuning API, which does not support image-based DPO, applying only BEAT w/o CTL on GPT-4o.
 Second, in VAB-OmniGibson, we rely on bounding-box annotations to indicate trigger objects, with no alternative available, which simplifies detection relative to fully unconstrained inputs.
 Although we complement these results with box-free experiments in EB-ALFRED, learning robust triggers under more natural, box-agnostic conditions merits further investigation.
+Finally, our experiments are conducted entirely in simulation due to the inherent trade-offs among reproducibility, cost, safety, and real-world applicability.
+Although the BEAT finetuning framework is not tied to any specific simulator, extending our evaluation to real-world environments remains an important direction for future work.
+Appendix G Ethics statement
+By exposing how visual backdoors can subvert embodied agents, our work serves as a double-edged sword: it highlights serious vulnerabilities that could be weaponized, but we believe transparency is essential to drive the development of effective defenses.
+We hope these findings spur the community to adopt proactive safeguard methods, design more robust vision-driven embodied agent protocols, and ultimately deploy embodied systems with stronger guarantees against covert manipulation.
+Appendix H Prompts
+H.1 Prompt for VAB-OmniGibson
+H.2 Prompt for EB-ALFRED